OpenStudy (anonymous):

How to do this ? If an arithmetic progression consists of 31 terms , in which the 16th term is m,then find sum of all the terms of this ap. ( 1 mark )

OpenStudy (anonymous):

\[m=a+15d \\ S_{31}=\frac{31}{2}(2a+30d)=\frac{31}{2}2(a+15)=31m\]

OpenStudy (anonymous):

thanks a lot @Jonask

OpenStudy (anonymous):

there is a mistake i.e \[\frac{ 31 }{ 2 }2(a+15d)=31(a+15d)=31m\]

OpenStudy (anonymous):

hmmmm! that's understandable @Salmon

OpenStudy (anonymous):

Ya..and I liked the way you explained it...

OpenStudy (anonymous):

Ya..and I liked the way you explained it...@Jonask

OpenStudy (anonymous):

I am doubtful. sum of terms from 1 to n, (1 +2 + 3...), is (n/2)(n+1), so 1 to 10 sum to 55, for example. This seems at odds with solution given..

OpenStudy (anonymous):

here we are not dealing with consecutive terms so we cant use the analogy of n/2(n+1),the numbers have a difference we didnt compute,we dont need to even comute it.....suppose \[d=a=1,m=16\\1+2+3+4...+31\\S_{31}=16\times 31=\frac{31}{2}(31+1)\] wich is of the form n/2(n+1)...

OpenStudy (anonymous):

An arithmetic progression increases a constant amount A between terms, thus you can turn it into a progression that looks like A(1+2+3...n) and thus you can use the equation I gave, (n/2)(N+1), multiplying what you get by A. Yes, some manipulation is required, but they ask us not to give the answers.